OpenStudy (anonymous):

18 is what percent of 67?

OpenStudy (jdoe0001):

18 is x% of 67 so 67 would be the 100% then... so what's "x"? well \(\large \begin{array}{ccllll} amount&\% \\\hline\\ 67&100\\ 18&x \end{array}\implies \cfrac{67}{18}=\cfrac{100}{x}\)

OpenStudy (anonymous):

I thought x/100=67/18

OpenStudy (jdoe0001):

you'd just put the amount and % lined up vertically

OpenStudy (anonymous):

I checked your answer it's right thanks

OpenStudy (jdoe0001):

yw